# Exploiting Vocabulary Frequency Imbalance in Language Model Pre-training

### v1

@@ 3.5 Skew-driven effects persist across corpora quality levels
To test whether the effects of increasing vocabulary size hold across datasets of varying quality, we conduct experiments on both FineWeb-Edu Penedo et al. (2024) and OpenWebText Gokaslan et al. (2019). Figure 1(a) and 1(b) demonstrate that the impact of increasing vocabulary size has a similar effect on both high-quality datasets (e.g., FineWeb-Edu Penedo et al. (2024)) and the lower-quality ones (e.g., OpenWebText Gokaslan et al. (2019)). Figure 1(c) further shows that the frequent $2,500$ words coincide with nearly 7575% each other, highlighting the universality of frequent vocabulary across different corpora. Overall, the findings indicate that vocabulary size expansion produces the same effects irrespective of corpus quality and that high-frequency words substantially overlap across datasets.
3.6 Global loss declines as lower complexity and sharper skew cut frequent-token uncertainty
@@
The effect of tokenized text complexity and relative token-frequency imbalance on language models has not yet been explored. Each prediction in a language model depends on previous tokens, and the conditional nature of these probabilities adds significant analytical complexity. However, it is possible to hypothesize that loss of rare token prediction weighs heavily in the overall loss, as conditional probability is often orders of magnitude lower and therefore yields much higher per-token losses Pinto et al. (2024); Mircea et al. (2024). To address this presumption, we empirically examine how reducing loss on frequent word prediction at the expense of rare word predictions affects overall model performance. Figure 2(a) shows that enlarging the vocabulary steadily lowers the average per-word loss for the frequent words, where the gap between $24\mathrm{K}$ and $196\mathrm{K}$ is roughly $0.1$ nats for the frequent $2,500$ words. Moreover, frequent words dominate the cross-entropy loss: frequent $2,500$ words contribute almost 8080% of the total loss regardless of vocabulary size. However, as the vocabulary expands, the loss associated with infrequent words also rises, reflecting their diminished conditional probabilities and degraded predictability on those tokens. Despite this penalty, Figure 2(b) reports an overall decline in global cross-entropy loss from roughly $3.179$ to $3.136$ nats as the vocabulary size grows, indicating that the benefits of reducing loss on frequent words prediction outweigh the drawbacks of poorer infrequent word estimates. These observations highlight a key takeaway: lower tokenized text complexity makes the model predict frequent words with less uncertainty, and since the overall objective is governed by their loss contribution, further skewness in the relative token-frequency distribution generally reduces the data complexity and cross-entropy loss even though it increases rare words’ losses. This finding is consistent with the existing study Rajaraman et al. (2024) on unigram models, explaining the benefits of increasing vocabulary size. We observe the same qualitative pattern in the OpenWebText (Appendix C).
(a) Frequent word loss on a 10B token FineWeb-Edu
(b) Global cross-entropy on a 10B token FineWeb-Edu
@@
Figure 2: Figure 2(a) illustrates that models with a larger vocabulary size reduce loss on the most frequent $2,500$ words in the FineWeb-Edu. It also reveals that frequent words account for nearly 8080% of the total loss, while loss on infrequent words grows with vocabulary size as their conditional probabilities fall due to data sparsity. Nevertheless, Figure 2(b) shows that the global cross-entropy loss declines as vocabulary size increases, demonstrating that the gains from lower loss on frequent words outweigh the losses from poorer infrequent word estimates.
@@ 4 Analysis
Although our results show that minimizing loss on high frequency tokens is crucial for reducing global cross-entropy, several questions remain unresolved. To pinpoint the causal chain from tokenizer design to model behavior, we now organize our analysis around three additional guiding questions.
@@
    Norm-Constraining and token-frequency Imbalance
    If we constrain input/output embedding norms to eliminate the impact of token-frequency imbalance, does cross-entropy loss increase4.1)?
  2. 2.
@@
The next three subsections answer newly added Q11–Q33 in turn.
(a) Frequent word loss gap of base and embednorm
(b) Global cross-entropy gap of base and embednorm
Figure 3: Constraining the norms of both input and output token embeddings to mitigate the impact of token-frequency imbalance during training increases average per-word loss on $2,500$ frequent words in the FineWeb-Edu as illustrated in figure 3(a), which in turn drives up the global cross-entropy loss on the FineWeb-Edu (figure 3(b)).
4.1 Constraining embedding norms increases cross-entropy
Higher token-frequency imbalance provides plentiful training examples for frequent words but far fewer for rare ones. Section 3.6 demonstrates that this reduces loss on frequent words while inflating it on rare words. We evaluate whether reducing the impact of token-frequency imbalance during pre-training leads to a performance degradation.
In language models, the token embedding norm of input and output embeddings reflects the token-frequency imbalance in training data. For input embeddings, every time a token shows up in a training batch, its token embedding receives gradient updates, and tokens that do not appear only undergo weight decay, causing their embedding norms to shrink over time Land and Bartolo (2024); Chung et al. (2024). In the output embedding layer, the target token in the output embeddings receives a stronger gradient than non target token embedding during training Mircea et al. (2024) 111we assume input and output embeddings are untied from each other (see the Appendix D for a detail explanation). Thus embedding norm of a frequent token is typically larger in both input and output embedding. Even though language models with pre-LN Xiong et al. (2020) normalize the final layer hidden states, frequent tokens still develop larger output-embedding norms and highly align with the final layer hidden states, inflating their logits and predicted probabilities and ultimately reducing their average per-word loss.
To reduce the impact of token-frequency imbalance during pre-training, we constrain the token embeddings of the input and output embedding layer to a unit norm, since the token embeddings of our base model already average unit length. Figure 3(a) demonstrates that constraining the norms of both input and output token embeddings to address token-frequency imbalance during training leads to a higher average loss for $2,500$ frequent words. This increases the global cross-entropy on the training dataset and lowers downstream performance as illustrated in figure 3(b) and figure 4(c) respectively. This observation demonstrates that placing equal emphasis on rare and frequent tokens undermines performance; ensuring minimal uncertainty when predicting frequent tokens is essential.
@@ 4.2 Frequent-word overlap explains downstream performance transfer
Reducing loss for frequent words is key to achieving a lower global cross-entropy loss. But does this effect carry over to downstream task accuracy? Although a strong link between lower loss and better downstream performance has been studied Huang et al. (2025); Kim et al. (2025), its rationale has not been investigated. To shed light on this phenomenon, we analyze the overlap of frequent words between the pre-training data and evaluation benchmarks and demonstrate that scaling the vocabulary size reduces cross-entropy loss during pre-training and on downstream tasks.
[…]
@@
Figure 4: Figure 4(a) demonstrates that the most frequent $2,500$ words in the FineWeb-Edu comprise nearly 727872-78% of the tokens in other downstream benchmark datasets as well as the $C$-Main-2023402023-40 Huang et al. (2024). ARC refers to ARC-Easy, and HS refers to HellaSwag. Figure 4(b) illustrates that a larger vocabulary reduces average per-word loss on frequent FineWeb-Edu words within the CC dataset, and demonstrates how this translates into lower global cross-entropy loss on CC dataset. Figure 4(c) confirms that scaling the vocabulary size boosts downstream task performance.
@@ 4.3 Parameter scaling recovers the same frequent-word gains
Thus far, we have shown that enlarging the vocabulary size decreases word-level average per-word loss on frequent words, which translates into lower cross-entropy loss and better downstream task performance. One might ask whether similar gains could be achieved without altering the vocabulary size by adjusting other model hyperparameters. Our experiment illustrates that increasing the model’s parameter count can replicate the same benefit. To investigate this, we use the Pythia suite, where each checkpoint is trained on the same dataset with identical hyperparameters, differing only in model size Biderman et al. (2023). We also examine the OLMo-2 series OLMo et al. (2025) to determine whether the same pattern persists in contemporary large-scale language models (see the Appendix E).
Figure 5 compares Pythia models (160160M, 11B, 6.96.9B parameter count) in terms of word-level average per-word loss, proportion of frequent words losses on total loss, and global cross-entropy loss measured on the Paloma validation dataset Magnusson et al. (2024). Figure 5(a) shows that larger models predict frequent words far more accurately: for the most frequent $2,500$ words in FineWeb-Edu, the average per-word loss drops from $4.48$ nats (160160M) to $3.76$ nats (11B) and $3.38$ nats (6.96.9B). Furthermore, frequent words loss dominates the total loss as model size grows, which accounts for almost 7070% of the total loss in the Pythia 11B and 6.96.9B model size, compared with roughly 4848% in the Pythia 160160M. Unlike the trend observed when only the vocabulary size is increased (Figure 2(b)), scaling model capacity does not inflate loss on rare tokens; their share of the loss shrinks. Figure 5(b) illustrates that the global cross-entropy loss on the Paloma validation set drops from $4.32$ nats in the Pythia 160160M to $2.51$ nats and $2.26$ nats in the Pythia 11B and 6.96.9B respectively, mirroring the loss reduction seen when expanding the vocabulary (Figure 3(b)). From this experiment, we can observe that enlarging the model size also lowers loss on frequent words while effectively ignoring rare token losses, so the global cross-entropy reduction is larger than what scaling up vocabulary size alone can deliver.
@@ 5 Discussion
Table 2: Upper bound of Kolmogorov complexity ($K(X^{N})$) and NCR for the 45.7 billion-byte FineWeb-Edu corpus tokenized with various SuperBPE variants. $t$ and Avg denote the stage-switch vocabulary threshold and an average downstream performance reported by (Liu et al., 2025), respectively. The SuperBPE variant with the highest average performance exhibits the lowest tokenized text complexity.
superBPE $K(X^{N})$ NCR Avg (Liu et al., 2025)
$200\mathrm{K}(BPE)$ $10.21B$ 0.223 39.8
$200\mathrm{K}(t=180K)$ $10.03B$ 0.219 43.8
$200\mathrm{K}(t=160K)$ $10.05B$ 0.220 43.4
$200\mathrm{K}(t=80K)$ $10.10B$ 0.221 42.9
@@ 5.1 Can we reduce the tokenized text complexity without intensifying frequency imbalance?
A larger vocabulary reduces the tokenized text complexity by decreasing the number of tokens and sharpening the token-frequency imbalance. However, section 3.6 shows that a sharper token-frequency imbalance increases rare-token loss. This raises the question: can we lower tokenized text complexity without enlarging the vocabulary size? SuperBPE (Liu et al., 2025) is one such example.
@@
SuperBPE employs a two-stage BPE algorithm: in the first stage (up to a threshold vocabulary size $t$), it operates identically to standard BPE, while in the second stage, it abandons whitespace pre-tokenization. By permitting merges across former whitespace boundaries, Superbpe directs subsequent merges toward frequent tokens, thereby limiting the introduction of new rare tokens as the vocabulary expands, and preventing further decrease of relative token frequency of rare tokens. Simultaneously, it markedly decreases the token count of data, yielding up to $33\%$ fewer tokens than standard BPE on average.
@@
Table 2 reports the upper bound Kolmogorov complexity, NCR of the 45.7 billion-byte FineWeb-Edu corpus tokenized with various SuperBPE variants, along with average downstream performance reported by (Liu et al., 2025). The result shows that the superBPE achieves a lower token count and a less skewed token-frequency distribution than BPE at equal vocabulary. Also, the superBPE variant with the highest average performance exhibits the lowest tokenized text complexity. This finding demonstrates that both vocabulary expansion and the superBPE variants leverage the same benefit: they reduce the complexity of the tokenized text by segmenting common character sequences into single tokens, thereby facilitating the model to learn non-i.i.d. patterns in the data more easily.
@@
5.2 Why higher token-frequency imbalance degrades machine translation task performance?
@@
Zouhar et al. (2023) reported that higher token-frequency imbalance degrades machine translation task performance, contrary to our findings. What explains this contrasting behavior between machine translation and monolingual settings? We presume that this is related to rare word issues in machine translation tasks Sennrich et al. (2016b); Sennrich and Zhang (2019); Koehn and Knowles (2017); Luong et al. (2015); Gu et al. (2020).
@@
In machine translation, the source and target vocabularies overlap minimally: vocabularies that are common in English often occur rarely in French, and vice versa. Conversely, in a monolingual setting, the pre-training corpus and downstream benchmarks share not only vocabularies but also many of the same frequent words, resulting in extensive text overlap. Consequently, BLEU penalizes every missing target-side n-gram, so mistranslating even a handful of infrequent tokens can sharply lower scores. To address this rare-word challenge, methods like vocabulary trimming deliberately shrink the vocabulary size to reduce the impact of low-frequency tokens on translation performance Gowda and May (2020); Provilkov et al. (2020); Chizhov et al. (2024).

### v2

@@ 3.5 Skew-driven effects persist across corpora quality levels
To test whether the effects of increasing vocabulary size hold across datasets of varying quality, we conduct experiments on both FineWeb-Edu [40] and OpenWebText [13]. Figure 1(a) and 1(b) demonstrate that the impact of increasing vocabulary size has a similar effect on both high-quality datasets (e.g., FineWeb-Edu [40]) and the lower-quality ones (e.g., OpenWebText [13]). Figure 1(c) further shows that the frequent $2,500$ words coincide with nearly 7575% each other, highlighting the universality of frequent vocabulary across different corpora. Overall, the findings indicate that vocabulary size expansion produces the same effects irrespective of corpus quality and that high‑frequency words substantially overlap across datasets.
3.6 Larger vocabularies reduce frequent words uncertainty and global loss
@@
The effect of tokenized text complexity and relative token-frequency imbalance on language models has not yet been explored. Each prediction in a language model depends on previous tokens, and the conditional nature of these probabilities adds significant analytical complexity. However, it is possible to hypothesize that loss of rare token prediction weighs heavily in the overall loss, as conditional probability is often orders of magnitude lower and therefore yields much higher per‑token losses [42, 36]. To address this presumption, we empirically examine how reducing loss on frequent word prediction at the expense of rare word predictions affects overall model performance. Models are pre-trained on $40\mathrm{B}$ bytes and evaluated on a separate, non-overlapping $5\mathrm{B}$ byte split of FineWeb-Edu. Figure 2(a) shows that enlarging the vocabulary steadily lowers the average per‑word loss for the frequent words, where the gap between $24\mathrm{K}$ and $196\mathrm{K}$ is roughly $0.1$ nats for the frequent $2,500$ words. In contrast, the average per-word loss for the rarest $20,000$ words rises with vocabulary size, increasing from roughly $11.183$ to $13.399$ nats. As shown in Figure 2(b), the global cross-entropy declines from about $3.179$ to $3.136$ nats, implying that gains on frequent-word prediction outweigh the degradation on infrequent words. Frequent tokens dominate the objective in all settings: the top 2,5002{,}500 words account for nearly $75\%$ of total loss. Nevertheless, as the vocabulary expands, losses on infrequent tokens grow, reflecting lower conditional probabilities and reduced predictability for those items. These observations highlight a key takeaway: once loss on frequent words declines, the overall objective is governed by their loss contribution, so further skewness in the token frequency distribution generally improves performance even though it harms rare-token estimates. This finding is consistent with the existing study [44, 35], explaining the benefits of increasing vocabulary size. We observe the same qualitative pattern in the OpenWebText (Appendix C).
(a) Frequent vs. rare word loss (85M, 10B)
(b) CE-loss & frequent-word coverage (85M, 10B)
@@
Figure 2: Figure 2(a) illustrates that models with a larger vocabulary size reduce loss on the most frequent $2,500$ words while increase loss on the rarest $20,000$ words. Nevertheless, Figure 2(b) shows that the global cross‐entropy loss declines as vocabulary size increases, demonstrating that the gains from lower loss on frequent words outweigh the losses from poorer infrequent word estimates. It further reveals that frequent words account for nearly 7575% of the total loss, while loss on infrequent words grows with vocabulary size as their conditional probabilities fall due to data sparsity. Models are pre-trained on $40\mathrm{B}$ bytes and evaluated on a disjoint $5\mathrm{B}$ byte split of FineWeb-Edu.
3.7 Gains from vocabulary expansion hold for larger datasets and models
To assess whether vocabulary size effects persist across data and model scales, we train an $85\mathrm{M}$ (non-embedding) model on a $30\mathrm{B}$ GPT-2 token subset ($184.6\mathrm{B}$ bytes) of FineWeb-edu and a $450\mathrm{M}$ (non-embedding) model on a $10\mathrm{B}$ token subset. On the $30\mathrm{B}$ token split, figure 3(a) shows frequent-word loss drops from $3.845$ to $3.742$ nats, while rare-word loss rises from $10.199$ to $11.787$ nats; figure 3(b) shows global cross-entropy falls from $2.991$ to $2.941$ nats and frequent-word loss coverage declines from $77.6\%$ to $74.2\%$, exhibit the same trend reported in Section 3.6. Figure 3(c) and 3(d) confirm the same pattern for the 450M models: frequent-word loss decreases from $3.770$ to $3.675$ nats, rare-word loss increases from $9.694$ to $12.762$ nats, cross-entropy reduces from $2.989$ to $2.888$ nats, and frequent word loss coverage from $78.0\%$ to $73.6\%$. Overall, these findings indicate that vocabulary expansion produces the same effects with larger dataset and model size.
(a) Frequent vs. rare word loss (85M, 30B)
(b) CE-loss & frequent-word coverage (85M, 30B)
(c) Frequent vs. rare word loss (450M, 10B)
(d) CE-loss & frequent-word coverage (450M, 10B)
Figure 3: For an $85\mathrm{M}$ model trained on $30\mathrm{B}$ tokens, larger vocabularies reduce the most frequent $2,500$ word loss while increase the rarest $20,000$ word loss; since frequent words dominate, global cross-entropy drops (figure 3(a) and 3(b)). $450\mathrm{M}$ model trained on 10B tokens mirrors the pattern (figure 3(c) and 3(d)), indicating that these vocabulary-size effects persist across larger datasets and models.
@@ 4 Analysis
Although our results show that minimizing loss on high frequency tokens is crucial for reducing global cross-entropy, several questions remain unresolved. To pinpoint the causal chain from tokenizer design to model behavior, we now organize our analysis around two additional guiding questions.
@@
The next two subsections answer newly added Q11–Q22 in turn.
@@ 4.1 Frequent-word overlap explains downstream performance transfer
Reducing loss for frequent words is key to achieving a lower global cross-entropy loss. But does this effect carry over to downstream task accuracy? Although a strong link between lower loss and better downstream performance has been studied [20, 24], its rationale has not been investigated. To shed light on this phenomenon, we analyze the overlap of frequent words between the pre-training data and evaluation benchmarks and demonstrate that scaling the vocabulary size reduces cross-entropy loss during pre-training and on downstream tasks. Figure 4(a) shows that the $2,500$ most common words in FineWeb-Edu [40] account for roughly 7676-7878% of all tokens in ARC [3], HellaSwag [63], and SciQ [58] and about 7272% in the PIQA [5] and CC‑Main‑2023402023‑40 dump [21]. Because lower cross‑entropy loss on this CC subset correlates closely with stronger reasoning benchmark scores [21], reducing the frequent words loss and driving down global cross-entropy will improve accuracy across downstream tasks. Consistent with that expectation, figure 4(b) shows that increasing the vocabulary size from $24\mathrm{K}$ to $196\mathrm{K}$ lowers the average per-word loss on the frequent $2,500$ FineWeb-Edu words by about $0.11$ nats, translating into a roughly $0.07$ nats drop in global cross-entropy loss on the CC dataset. Figure 4(c) further confirms that a larger vocabulary also improves average downstream task accuracy of the language model across datasets and model scales. The key insight from this observation is that frequent words highly overlap between the typical training dataset and the downstream benchmark, so the cross‑entropy reduction achieved by enlarging the vocabulary size during pre‑training naturally translates into better downstream performance.
@@
Figure 4: Figure 4(a) demonstrates that the most frequent $2,500$ words in the FineWeb-Edu comprise nearly 727872-78% of the tokens in other downstream benchmark datasets as well as the $C$-Main-2023402023-40 [21]. ARC refers to ARC-Easy, and HS refers to HellaSwag. Figure 4(b) illustrates that a larger vocabulary reduces average per-word loss on frequent FineWeb-Edu words within the CC dataset, and demonstrates how this translates into lower global cross-entropy loss on CC dataset. Figure 4(c) confirms that scaling the vocabulary size boosts downstream task performance.
@@ 4.2 Parameter scaling recovers the same frequent-word gains
Thus far, we have shown that enlarging the vocabulary size decreases word-level average per-word loss on frequent words, which translates into lower cross‑entropy loss and better downstream task performance. One might ask whether similar gains could be achieved without altering the vocabulary size by adjusting other model hyperparameters. Our experiment illustrates that increasing the model’s parameter count can replicate the same benefit. To investigate this, we use the Pythia suite, where each checkpoint is trained on the same dataset with identical hyperparameters, differing only in model size [4]. We also examine the OLMo-2 series [38] to determine whether the same pattern persists in contemporary large-scale language models (see the Appendix D).
Figure 5 compares Pythia models (160160M, 11B, 6.96.9B parameter count) in terms of word-level average per-word loss, proportion of frequent words losses on total loss, and global cross-entropy loss measured on the Paloma validation dataset [35]. Figure 5(a) shows that larger models predict frequent words far more accurately: for the most frequent $2,500$ words in FineWeb-Edu, the average per‑word loss drops from $4.48$ nats (160160M) to $3.76$ nats (11B) and $3.38$ nats (6.96.9B). Furthermore, frequent words loss dominates the total loss as model size grows, which accounts for almost 7070% of the total loss in the Pythia 11B and 6.96.9B model size, compared with roughly 4848% in the Pythia 160160M. Unlike the trend observed when only the vocabulary size is increased (Figure 2(b)), scaling model capacity does not inflate loss on rare tokens; their share of the loss shrinks. Figure 5(b) illustrates that the global cross-entropy loss on the Paloma validation set drops from $4.32$ nats in the Pythia 160160M to $2.51$ nats and $2.26$ nats in the Pythia 11B and 6.96.9B respectively, mirroring the loss reduction seen when expanding the vocabulary (Figure LABEL:fig:embd_b). From this experiment, we can observe that enlarging the model size also lowers loss on frequent words while effectively ignoring rare token losses, so the global cross‑entropy reduction is larger than what scaling up vocabulary size alone can deliver.
@@ 5 Discussion
Table 2: Upper bound of Kolmogorov complexity ($K(X^{N})$) and NCR for the $45.97$ billion‑byte FineWeb‑Edu corpus tokenized with various SuperBPE variants. $t$ and Avg denote the stage-switch vocabulary threshold and an average downstream performance reported by [32], respectively. The SuperBPE variant with the highest average performance exhibits the lowest tokenized text complexity.
superBPE $K(X^{N})$ NCR Avg [32]
$200\mathrm{K}(BPE)$ $10.21B$ 0.222 39.8
$200\mathrm{K}(t=180K)$ $10.03B$ 0.218 43.8
$200\mathrm{K}(t=160K)$ $10.05B$ 0.219 43.4
$200\mathrm{K}(t=80K)$ $10.10B$ 0.220 42.9
@@ 5.1 Can we reduce the tokenized text complexity without intensifying frequency imbalance?
A larger vocabulary reduces the tokenized text complexity by decreasing the number of tokens and sharpening the token-frequency imbalance. However, Section 3.6 shows that a sharper token-frequency imbalance increases rare-token loss. This raises the question: can we lower tokenized text complexity without enlarging the vocabulary size? SuperBPE [32] is one such example.
@@
SuperBPE employs a two‑stage BPE algorithm: in the first stage (up to a threshold vocabulary size $t$), it operates identically to standard BPE, while in the second stage, it abandons whitespace pre‑tokenization. By permitting merges across whitespace boundaries, Superbpe directs subsequent merges toward frequent tokens, thereby limiting the introduction of new rare tokens, and preventing further token frequency imbalance as the vocabulary expands.
@@
Table 2 reports the upper bound Kolmogorov complexity, NCR of the $45.97$ billion‑byte FineWeb‑Edu corpus tokenized with various SuperBPE variants, along with average downstream performance reported by [32]. The result shows that the superBPE achieves a lower token count and a less skewed token-frequency distribution than BPE at equal vocabulary. Also, the superBPE variant with the highest average performance exhibits the lowest tokenized text complexity. This finding demonstrates that both vocabulary expansion and the superBPE variants leverage the same benefit: they reduce the complexity of the tokenized text by segmenting common character sequences into single tokens, thereby facilitating the model to learn non-i.i.d. patterns in the data more easily.
5.2 Deduplication through the lens of vocabulary frequency imbalance
Deduplication removes exact or near-duplicate content from a corpus, reducing repetition and leakage so that a larger fraction of the dataset is novel [28, 40, 31]. Deduplication can increase the information density of a dataset, i.e., the fraction of non-redundant data in the corpus. For non-i.i.d. data, information density can be proxied by the byte-level entropy rate, the theoretical minimum amount of non-redundant information. However, the entropy rate is only tightly approximable with near infinite context compressors, finite window compressors can distort comparisons before and after deduplication. A more practical loose upper bound is the byte-level bits-per-byte (BPB): under an i.i.d. assumption it quantifies information density, with unigram (byte) entropy upper-bounding the true entropy rate. Under the byte-level vocabulary, greater information density implies a higher BPB, driving entropy toward $\log_{2}256=8$ bits/byte, which corresponds to a more uniform distribution and lower frequency imbalance. Exploring how deduplication alters BPB, information density, and vocabulary frequency imbalance is a promising avenue for future work.
@@ 5.2 Deduplication through the lens of vocabulary frequency imbalance
5.3 Why higher token‑frequency imbalance degrades machine translation task performance?
@@
Zouhar et al. [64] reported that higher token-frequency imbalance degrades machine translation task performance, contrary to our findings. What explains this contrasting behavior between machine translation and monolingual settings? We presume that this is related to rare word issues in machine translation tasks [52, 50, 25, 34, 17].
@@
In machine translation, the source and target vocabularies overlap minimally: vocabularies that are common in English often occur rarely in French, and vice versa. Conversely, in a monolingual setting, the pre-training corpus and downstream benchmarks share not only vocabularies but also many of the same frequent words, resulting in extensive text overlap. Consequently, BLEU penalizes every missing target-side n-gram, so mistranslating even a handful of infrequent tokens can sharply lower scores. To address this rare-word challenge, methods like vocabulary trimming deliberately shrink the vocabulary size to reduce the impact of low-frequency tokens on translation performance [15, 43, 8].